how do you do 2.85y-7=12.85y-2 step by step
Found 2 solutions by solver91311, edjones:
Answer by solver91311(17077)



Add -2.85y to both sides.

Add 2 to both sides.

Multiply both sides by 1/10.


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by edjones(7569)

2.85y-7=12.85y-2
-10y=5
y=-1/2
.
Ed